Cynthia Besch wants to buy a rug for a room that is 18 ft wide and 19 ft long. She wants to leave a uniform strip of floor around the rug. She can afford to buy 210 square feet of carpeting. What dimensions should the rug​ have?
Answer by jorel1380(3719)

Let w be the width of the strip. Then the dimensions of the rug would be 19-2w and 18-2w. So:
(19-2w)(18-2w)=210
342-74w+4w²=210
4w²-74w+132=0
(2w-33)(2w-4)=0
w=33/2 or 2
A width of 2 ft for the strip gives rug dimensions of 15 ft. x 14 ft. ☺☺☺☺